# GenView: Enhancing View Quality with Pretrained Generative Model for Self-Supervised Learning

### v1

@@ Appendix A Implementation Details
Comparison with other view construction methods. We compare our approach with several baseline methods, including ContrastiveCrop [65] (C-Crop), ViewMaker [80], Neural Transform Network [42] (NTN), Local Manifold Augmentation method [95] (LMA), Diffusion-based augmentation from scratch [98] (DiffAug), and $\mathcal{W}$-perturb [30]. For our experiments, we use three datasets: CF10, CF100, and TinyIN. We employ SGD as the optimizer with a learning rate of 0.5, weight decay of 5e-4, and momentum of 0.9. The learning rate follows a linear warm-up for 10 epochs and then switches to the cosine decay scheduler. Batch sizes are set to 512 for CF10 and CF100 and 256 for TinyIN. The momentum coefficient for the momentum-updated encoder and memory buffer size is set to 0.99/0.99/0.996 and 4096/4096/16384 for CF10/CF100/TinyIN, respectively. We use the $\mathcal{L}_{\text{SSL, NCE}}$ loss with a temperature of 0.2 and train for 500 epochs on CF10 and CF100 and 200 epochs on TinyIN. The backbone architecture used is ResNet18 with an embedding dimension of 512 and a projection dimension of 128. We replace the first 7x7 Conv of stride 2 with 3x3 Conv of stride 1 and remove the first max-pooling operation. For data augmentations, we use random resized crops (the lower bound of random crop ratio is set to 0.2), color distortion (strength=0.4) with a probability of 0.8, and Gaussian blur with a probability of 0.5. The images from the CF10/CF100 and TinyIN datasets are resized to 32x32 and 64x64 resolution.
Appendix B Additional Illustration
Further visualization and limitation analysis. Fig. 5 displays positive pairs generated by GenView, showing its ability to introduce variations while maintaining semantic consistency. Notable observations include:
@@ Appendix B Additional Illustration
Evaluation of positive views constructed by different methods. We calculate the average cosine similarity between the original images and their associated positive views for different methods. We randomly sample 50,000 images from IN-1K and compute the cosine similarity of CLIP image embeddings for each pair. The results are presented in Tab. 10, which demonstrates that our method produces positive views with significantly higher semantic similarity to the original images compared to the RS and Retrieval Laion400M methods.
Appendix C Algorithm
@@
19:     $\mathcal{L}_{\text{SSL,*}}\leftarrow\text{ComputeContrastiveLoss}(\mathbf{P}_{i},\mathbf{P}_{i}^{+})$

### v2

@@ Appendix 0.B Additional Illustration
Fig. 5 displays positive pairs generated by GenView, showing its ability to introduce variations while maintaining semantic consistency. Notable observations include:
@@
Evaluation of positive views constructed by different methods.
We calculate the average cosine similarity between the original images and their associated positive views for different methods. We randomly sample 50,000 images from IN-1K and compute the cosine similarity of CLIP image embeddings for each pair. The results are presented in Tab. 10, which demonstrates that our method produces positive views with significantly higher semantic similarity to the original images compared to the RS and Retrieval Laion400M methods.
Appendix 0.C Algorithm
@@
19:     $\mathcal{L}_{\text{SSL,*}}\leftarrow\text{ContrastiveLoss}(\mathbf{P}_{i},\mathbf{P}_{i}^{+})$